A 200 kg load is hung on a wire of length 4.00 m, cross-sectional area 0.200 x 1024 m2 and Young’s modulus 8.00 x 1010N/m2. What is its increase in length?
1
Expert's answer
2020-10-27T08:23:14-0400

Young's Modulus(Y) is given by-

"Y=\frac{mgl}{ \pi r^2 \Delta l}"

Putting the values from the question-

"8\times10^{10}=\frac{(200)(9.81)(4)}{(0.2\times10^{-24})(\Delta l)}"

"\Delta l = 4.905\times10^{-14} m"